OpenStudy (anonymous):

Company A rents copiers for a monthly charge of $200 plus 8 cents per copy.Company B rents copiers for a monthly charge of $400 plus 4 cents per copy.What is the number of copies above which Company A's charges are the higher of the two?

OpenStudy (sgadi):

200+0.08*x>400+0.04*x 0.04*x>200 x>200/0.04 x>5000

OpenStudy (anonymous):

COULD YOU SIMPLIFY YOUR ANSWER a bit mr sgadi

OpenStudy (sgadi):

if company takes x copies per month, then cost for compant A is \[200+0.08x\] and cost for company B is \[200+0.04x\] so we need condition when cost of company A is greater or equal to second one .. so we can say 200+0.08*x>=400+0.04*x 0.04*x>=200 x>=200/0.04 x>=5000